# Learning from Suboptimal Data in Continuous Control via Auto-Regressive Soft Q-Network

### v1

@@ Abstract
Reinforcement Learning
@@ 1 Introduction
Deep reinforcement learning (RL) has demonstrated remarkable performance across various continuous control domains (Haarnoja et al., 2018; Schulman et al., 2017). However, these breakthroughs often come at the cost of extensive online interactions, which are required for effective convergence (Berner et al., 2019; Mnih et al., 2015). This reliance on large-scale exploration poses a major challenge in real-world applications, where data collection can be expensive, time-consuming, or even risky. To alleviate this burden, value-based RL methods, which directly approximate the Q-function rather than parameterizing a policy, have gained popularity due to their improved sample efficiency (Seyde et al., 2024; Tavakoli et al., 2020; Seyde et al., 2022) and have shown advances in continuous control tasks by discretizing each of the dimensions of continuous action spaces. (Seo et al., 2024). Moreover, some studies integrate offline demonstration data into training to further accelerate early learning, reducing the dependence on purely online exploration (Ball et al., 2023). In this paper, we adopt this training paradigm to address continuous control using value-based RL, incorporating offline data into the online training process.
For value-based RL, the discretization scheme results in an exponentially large discrete action space, making RL training and exploration challenging. To mitigate this, existing value-based methods often estimate the Q-value for each action dimension independently (Metz et al., 2017; Seyde et al., 2022). However, this simplification comes with a limitation—it neglects interdependencies between action dimensions, potentially leading to suboptimal decision-making. When training data exhibits multiple modes, such as a mix of optimal and suboptimal demonstrations, independently estimating Q-values can bias action selection toward the most frequent behaviors rather than the truly optimal ones. This limitation is particularly pronounced in the early stages of learning, when the agent relies heavily on imperfect offline data and lacks sufficient online refinement.
@@ 2 Related Works
Despite their inherently straightforward critic-only framework, value-based reinforcement learning (RL) algorithms have achieved notable success (Mnih et al., 2015; Silver et al., 2017; Schrittwieser et al., 2020; Seyde et al., 2024; Seo et al., 2024). Although these algorithms are primarily designed for discrete action spaces, recent efforts have sought to adapt them to continuous control by discretizing the continuous action space (Tavakoli et al., 2018; Seyde et al., 2022). However, the curse of dimensionality remains a significant challenge, as the number of discretization bins increases exponentially with the action dimension (Lillicrap, 2015). To address this issue, some studies have modified the Markov Decision Process (MDP) of the environment, transforming it into a sequential decision-making problem along the action dimension (Metz et al., 2017; Chebotar et al., 2023). Other approaches treat each action dimension independently, generating the Q function separately for each dimension (Tavakoli et al., 2018, 2020; Seyde et al., 2022, 2024), akin to treating each action dimension as a multi-agent RL problem (Foerster et al., 2018; Yu et al., 2022). Recent research (Seo et al., 2024) has employed a coarse-to-fine discretization approach to improve sample efficiency. However, treating each action dimension independently may disrupt the correlation between different action dimensions, potentially diminishing performance in policy optimization. Some studies (Seo & Abbeel, 2024) have attempted to solve this issue through action sequence prediction. Our approach generates actions in an auto-regressive manner, considering the correlations between dimensions and improving policy learning, which is orthogonal to (Seo & Abbeel, 2024).
@@
Deep reinforcement learning often requires a large amount of online interactions to achieve convergence (Berner et al., 2019; Mnih et al., 2015). To address this challenge, many methods have been proposed that leverage offline demonstrations to guide online exploration and accelerate policy training (Rajeswaran et al., 2017; Ball et al., 2023). Some approaches involve performing offline RL pretraining before initiating online RL training (Lee et al., 2022; Nakamoto et al., 2024; Lei et al., 2023; Hu et al., 2024). However, these approaches often depend on expensive offline pretraining. To mitigate this, some works explore incorporating offline demonstration data directly into the training process. One strategy initializes the replay buffer with offline data (Hester et al., 2018; Ball et al., 2023), while another balances sampling between online and offline data to improve training stability (Zhang et al., 2023; Hansen et al., 2022). Additionally, certain methods explicitly introduce a behavior cloning loss to leverage high-quality demonstrations for better guidance (Rudner et al., 2021; Rajeswaran et al., 2017; Nair et al., 2018). In this work, we adopt the paradigm of integrating offline demonstrations into training to enhance sample efficiency in continuous control tasks. Specifically, we improve value-based RL by introducing an auto-regressive structure that sequentially estimates advantage for each action dimension. This design enables better handling of suboptimal data, whether from offline demonstrations or trajectories collected during training.
@@ 3.2 Soft Q Learning
where $\mathcal{H}$ is entropy, $T$ is the episode length and $\rho_{\pi}$ is the trajectory distribution induced by policy $\pi$. The temperature parameter $\alpha$ dictates how much importance is placed on the entropy term in comparison to the reward. Let the soft Q-function defined as:
@@ 4.2 Dimensional Soft Advantage for Policy Representation
However, outputting such a large number of Q values imposes a significant computational burden on the neural network. Inspired by auto-regression (Brown et al., 2020), we address this problem by generating the Q function for a state-action pair in an auto-regressive manner.
For clarity, we treats discrete action discussed in Sec. 4.1 in one level. The multi-level coarse-to-fine discrete action can be considered as additional action dimensions, without compromising generalization. We first define the dimensional soft advantage for policy representation.
@@
The dimensional soft advantage of the action $a^{d}$ at state $\mathbf{s}$, considering the previous dimensional actions $\mathbf{a}^{-d}=(a^{1},\cdots,a^{d-1})$, is expressed by
$\pi(a^{d}|\mathbf{s},\mathbf{a}^{-d})=\frac{exp\left(\frac{1}{\alpha}A^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d}))\right)}{Z(\mathbf{s},\mathbf{a}^{-d})}$ (11)
where $Z^{d}(\mathbf{s},\mathbf{a}^{-d})$ represents the partition function.
However, the dimensional soft advantage is not related to other equations and remains intractable. To address this, we propose the following theorem to establish a connection between the dimensional soft advantage and the soft advantage.
@@
If dimensional soft advantage $m^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d}))$ satisfies
$log\sum_{a^{d^{\prime}}}\exp{\left(\frac{1}{\alpha}A^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d^{\prime}})\right)}=0$ (12)
then the soft advantage can then be expressed as the summation of the dimensional soft advantages
@@
Through Eq. (11) and Theorem 4.3, we extend soft Q-learning to a auto-regressive policy along action dimension.
@@ 4.3 Auto-Regressive Soft Q-learning
$-log\sum_{a^{d^{\prime}}}\exp\left(\frac{1}{\alpha}u^{d}(\mathbf{s}_{t},\mathbf{a}^{-d},a^{d^{\prime}})\right)$
@@
Furthermore, to stabilize training and address the over-estimation problem (Fujimoto et al., 2018; Van Hasselt et al., 2016), we implemented a double Q network alongside a target network in our practical application. Therefore, the optimization objective in Eq. (14) is modified to
$\mathbf{y}_{t}=\gamma\mathbb{E}_{\mathbf{s}_{t+1}\sim p(s)}\left[\text{min}\left(V^{\overline{\phi}_{1}}_{\text{soft}1}(\mathbf{s}_{t+1}),V^{\overline{\phi}_{2}}_{\text{soft}2}(\mathbf{s}_{t+1})\right)\right]$ (17)
$\mathcal{L}_{RL}=\frac{1}{2}\left(V^{\phi_{i}}_{\text{soft}}(\mathbf{s}_{t})+A^{\theta_{i}}(\mathbf{s}_{t},\mathbf{a}_{t})-\mathbf{y}_{t}\right)$ (18)
where $V^{\overline{\phi}_{i}}_{\text{soft}}$ represents the soft value predicted by the target network.
@@ 5 Experiment
Baselines. We use CQN (Seo et al., 2024), a state-of-the-art value-based RL method for continuous control, as our baseline. CQN employs a coarse-to-fine action selection strategy and independently predicts Q-values for each action dimension. Additionally, CQN trains using a combination of online training and offline demonstrations. Besides, we also include DrQ-v2 (Yarats et al., 2021), a renowned actor-critic algorithm designed for vision-based RL, along with its enhanced version, DrQ-v2+, as benchmarks. We also feature ACT (Zhao et al., 2023) and a CQN-style behavior cloning (BC) policy among our baselines. Details about the baselines can be found in Appendix C.3.
@@ 5.1 Performance on D4RL
To better investigate the influence of dataset quality, we rank trajectories by episode return for each dataset, and labeling the top $30$%, middle $30$%, and bottom $30$% of the data as offline demonstrations. The behavior cloning objective is applied only to these offline demonstrations. We report the converged performance ARSQ, CQN and BC. As illustrated in Fig. 5, ARSQ consistently outperforms both CQN and BC across all three levels of demonstration quality. Notably, when using the bottom $30$% of data as offline demonstrations, ARSQ achieves approximately $\times2$ the final performance of CQN. In contrast, with the lowest demonstration quality, CQN performs slightly worse than BC, revealing CQN’s sensitivity to demonstration quality, which negatively affects its online training. These results further validate the effectiveness of our method when using suboptimal offline datasets.
@@ 5.2 Performance on RLBench
In this domain, we evaluate the performance of ARSQ, CQN, DrQ-v2+, DrQ-v2, ACT and BC. All reinforcement learning methods incorporate the behavior cloning objective (Eq. (15)) towards expert demonstrations and successful trajectories collected online.
@@ 5.3 Ablation Studies
Ablation on auto-regressive conditioning
@@
As depicted in Fig. 7, Swap demonstrates a slight decline in performance, underscoring the effectiveness of the current conditioning order design. Additionally, removing any of the components degrades performance to varying degrees. When all components are removed, as in Plain, the performance is at its lowest, emphasizing the significance of dimensional and coarse-to-fine action generation.
@@
As shown in Fig. 8, the standard ARSQ consistently performs well in both environments. In contrast, using either a level-shared or separate backbone results in diminished performance. This demonstrates the effectiveness of the shared backbone design.
@@ B.3 Network Architecture
In RLBench tasks, observations consist of a combination of RGB images and low-dimensional states. To compute the dimensional soft advantage for a given dimension, we first input the RGB images and low-dimensional states into a Convolutional Neural Network (CNN) (Li et al., 2021) encoder and a Multi-Layer Perceptron (MLP) (Haykin, 1998) encoder, respectively, to extract feature representations. These representations are then used to predict the soft value. Concurrently, the feature representations are combined with actions from previous dimensions and coarse-to-fine levels to create auto-regressive conditioning. An MLP-based shared backbone and output head are then utilized to determine the dimensional soft advantage for the given dimension.

### v2

@@ Abstract
Value-based Reinforcement Learning, Continuous Control, Auto-Regressive, Suboptimal Demonstration
@@ 1 Introduction
Deep reinforcement learning (RL) has demonstrated remarkable performance across various continuous control domains (Haarnoja et al., 2018; Schulman et al., 2017). However, these breakthroughs often come at the cost of extensive online interactions, which are required for effective convergence (Berner et al., 2019; Mnih et al., 2015). This reliance on large-scale exploration poses a major challenge in real-world applications, where data collection can be expensive, time-consuming, or even risky. To alleviate this burden, value-based RL methods, which directly approximate the Q-function rather than parameterizing a policy, have gained popularity due to their improved sample efficiency (Seyde et al., 2024; Tavakoli et al., 2021; Seyde et al., 2023) and have shown advances in continuous control tasks by discretizing each of the dimensions of continuous action spaces (Seo et al., 2024). Moreover, some studies integrate offline demonstration data into training to further accelerate early learning, reducing the dependence on purely online exploration (Ball et al., 2023). In this paper, we adopt this training paradigm to address continuous control using value-based RL, incorporating offline data into the online training process. Project page is https://sites.google.com/view/ar-soft-q.
For value-based RL, the discretization scheme results in an exponentially large discrete action space, making RL training and exploration challenging. To mitigate this, existing value-based methods often estimate the Q-value for each action dimension independently (Metz et al., 2017; Seyde et al., 2023). However, this simplification comes with a limitation—it neglects interdependencies between action dimensions, potentially leading to suboptimal decision-making. When training data exhibits multiple modes, such as a mix of optimal and suboptimal demonstrations, independently estimating Q-values can bias action selection toward the most frequent behaviors rather than the truly optimal ones. This limitation is particularly pronounced in the early stages of learning, when the agent relies heavily on imperfect offline data and lacks sufficient online refinement.
@@ 2 Related Works
Despite their inherently straightforward critic-only framework, value-based reinforcement learning (RL) algorithms have achieved notable success (Mnih et al., 2015; Silver et al., 2017; Schrittwieser et al., 2020; Seyde et al., 2024; Seo et al., 2024). Although these algorithms are primarily designed for discrete action spaces, recent efforts have sought to adapt them to continuous control by discretizing the continuous action space (Tavakoli et al., 2018; Seyde et al., 2023). However, the curse of dimensionality remains a significant challenge, as the number of discretization bins increases exponentially with the action dimension (Lillicrap, 2015). To address this issue, some studies have modified the Markov Decision Process (MDP) of the environment, transforming it into a sequential decision-making problem along the action dimension (Metz et al., 2017; Chebotar et al., 2023). Other approaches treat each action dimension independently, generating the Q function separately for each dimension (Tavakoli et al., 2018, 2021; Seyde et al., 2023, 2024), akin to treating each action dimension as a multi-agent RL problem (Foerster et al., 2018; Yu et al., 2022). Recent research (Seo et al., 2024) has employed a coarse-to-fine discretization approach to improve sample efficiency. However, treating each action dimension independently may disrupt the correlation between different action dimensions, potentially diminishing performance in policy optimization. Some studies (Seo & Abbeel, 2024) have attempted to solve this issue through action sequence prediction. Our approach generates actions in an auto-regressive manner, considering the correlations between dimensions and improving policy learning, which is orthogonal to (Seo & Abbeel, 2024).
@@
Deep reinforcement learning often requires a large amount of online interactions to achieve convergence (Berner et al., 2019; Mnih et al., 2015). To address this challenge, many methods have been proposed that leverage offline demonstrations to guide online exploration and accelerate policy training (Rajeswaran et al., 2018; Ball et al., 2023). Some approaches involve performing offline RL pretraining before initiating online RL training (Lee et al., 2022; Nakamoto et al., 2023; LEI et al., 2024; Hu et al., 2024). However, these approaches often depend on expensive offline pretraining. To mitigate this, some works explore incorporating offline demonstration data directly into the training process. One strategy initializes the replay buffer with offline data (Hester et al., 2018; Ball et al., 2023), while another balances sampling between online and offline data to improve training stability (Zhang et al., 2023; Hansen et al., 2023). Additionally, certain methods explicitly introduce a behavior cloning loss to leverage high-quality demonstrations for better guidance (Rudner et al., 2021; Rajeswaran et al., 2018; Nair et al., 2018). In this work, we adopt the paradigm of integrating offline demonstrations into training to enhance sample efficiency in continuous control tasks. Specifically, we improve value-based RL by introducing an auto-regressive structure that sequentially estimates advantage for each action dimension. This design enables better handling of suboptimal data, whether from offline demonstrations or trajectories collected during training.
@@ 3.2 Soft Q Learning
where $\mathcal{H}$ is entropy, $T$ is the episode length and $\rho_{\pi}$ is the trajectory distribution induced by policy $\pi$. The temperature parameter $\alpha$ dictates how much importance is placed on the entropy term in comparison to the reward. Let the soft Q-function and soft value function defined as:
@@ 4.2 Dimensional Soft Advantage for Policy Representation
However, outputting such a large number of Q values imposes a significant computational burden on the neural network. Inspired by auto-regression (Brown et al., 2020), we address this problem by make the policy $\pi$ generate action $\mathbf{a}=(a^{1},a^{2},\ldots,a^{D})$ auto-regressively along the action dimensions.
For clarity, we treats discrete action discussed in Sec. 4.1 in one level. The multi-level coarse-to-fine discrete action can be considered as additional action dimensions, without compromising generalization. We first define the dimensional soft advantage to represent the auto-regressive policy at dimension $d$.
@@
The dimensional soft advantage of the action $a^{d}$ at state $\mathbf{s}$, considering the previous generated dimensional actions $\mathbf{a}^{-d}=(a^{1},\cdots,a^{d-1})$, is expressed by
$\pi(a^{d}|\mathbf{s},\mathbf{a}^{-d})\propto exp\left(\frac{1}{\alpha}A^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d}))\right)$ (11)
However, the dimensional soft advantage is not normalized for each action dimension $d$. We propose the following theorem to establish a connection between the dimensional soft advantage and the soft advantage.
@@
If the dimensional soft advantage $A^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d})$ satisfies
$\sum_{a^{d}}\exp{\left(\frac{1}{\alpha}A^{d}(\mathbf{s},\mathbf{a}^{-d},a^{d})\right)}=1$ (12)
for all dimension $d$, then the soft advantage can then be expressed as the summation of the dimensional soft advantages
@@
Additionally, Eq. (12) shows that the exponential of the dimensional soft advantage represent a valid probability distribution. Using Eq. (11) together with Theorem 4.3, the dimensional soft advantage serves as a bridge between policy representation and Q prediction.
@@ 4.3 Auto-Regressive Soft Q-learning
$-\alpha log\sum_{a^{d^{\prime}}}\exp\left(\frac{1}{\alpha}u^{d}(\mathbf{s}_{t},\mathbf{a}^{-d},a^{d^{\prime}})\right)$
@@
Furthermore, to stabilize training and address the over-estimation problem (Fujimoto et al., 2018; van Hasselt et al., 2016), we implement a double Q-learning approach with two separate value networks and their corresponding target networks. Specifically, we maintain two online value networks $V^{\phi_{1}}_{\text{soft}}$ and $V^{\phi_{2}}_{\text{soft}}$, along with their respective target networks $V^{\overline{\phi}_{1}}_{\text{soft}}$ and $V^{\overline{\phi}_{2}}_{\text{soft}}$. Each target network is updated as an exponential moving average (EMA) of its respective online network parameters. The value target is then computed by taking the minimum of the two target networks’ predictions
$\mathbf{y}_{t}=\gamma\mathbb{E}_{\mathbf{s}_{t+1}\sim p(s)}\left[\min\left(V^{\overline{\phi}_{1}}_{\text{soft}}(\mathbf{s}_{t+1}),V^{\overline{\phi}_{2}}_{\text{soft}}(\mathbf{s}_{t+1})\right)\right]$ (17)
Thus the resulting optimization objective becomes
$\mathcal{L}_{RL}=\frac{1}{2}\left(V^{\phi_{i}}_{\text{soft}}(\mathbf{s}_{t})+A^{\theta_{i}}(\mathbf{s}_{t},\mathbf{a}_{t})-\mathbf{y}_{t}\right)^{2}$ (18)
where $A^{\theta_{i}}$ is the soft advantage function parameterized by $\theta_{i}$.
@@ 5 Experiment
Baselines. We use CQN  (Seo et al., 2024), a state-of-the-art value-based RL method for continuous control, as our baseline. CQN employs a coarse-to-fine action selection strategy and independently predicts Q-values for each action dimension. Additionally, CQN trains using a combination of online training and offline demonstrations. Besides, we also include DrQ-v2  (Yarats et al., 2022), a renowned actor-critic algorithm designed for vision-based RL, along with its enhanced version, DrQ-v2+, as benchmarks. We also feature ACT  (Zhao et al., 2023) and a CQN-style behavior cloning (BC) policy among our baselines. Details about the baselines can be found in Appendix C.3.
@@ 5.1 Performance on D4RL
To better investigate the influence of dataset quality, we rank trajectories by episode return for each dataset, and labeling the top $30$%, middle $30$%, and bottom $30$% of the data as offline demonstrations. The behavior cloning objective is applied only to these offline demonstrations. We report the converged performance ARSQ, CQN and BC over three random seeds. As illustrated in Fig. 5, ARSQ consistently outperforms both CQN and BC across all three levels of demonstration quality. Notably, when using the bottom $30$% of data as offline demonstrations, ARSQ achieves approximately 2.0×2.0$\times$2.0 × the final performance of CQN. In contrast, with the lowest demonstration quality, CQN performs slightly worse than BC, revealing CQN’s sensitivity to demonstration quality, which negatively affects its online training. These results further validate the effectiveness of our method when using suboptimal offline datasets.
@@ 5.2 Performance on RLBench
In this domain, we evaluate the performance of ARSQ, CQN, DrQ-v2+, DrQ-v2, ACT and BC. All reinforcement learning methods incorporate the behavior cloning objective (Eq. (15)) towards expert demonstrations and successful trajectories collected online. Results are averaged over three random seeds.
@@
5.3 Performance under Fully Offline Setting
To further examine the performance of ARSQ, we conduct experiments in a fully offline setting, where the algorithm learns solely from a predetermined dataset. We utilize nine locomotion datasets from D4RL, as outlined in Sec. 5.1. For offline RL methods, we employ CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022), TD3+BC (Fujimoto & Gu, 2021), Onestep RL (Brandfonbrener et al., 2021), and RvS-R (Emmons et al., 2022) as baselines. For offline imitation learning methods capable of handling suboptimal data, we use filtered BC  (Chen et al., 2021; Emmons et al., 2022), Decision Transformer (Chen et al., 2021), and DWBC (Xu et al., 2022) as baselines. For DWBC, We adopt its best performance under “Setting 2” from its original paper (Xu et al., 2022), which mark top 5% trajectories as expert trajectories based on total reward, and we re-evaluate DWBC under the same conditions on additional datasets.
Table 1: Performance under fully offline setting.
Dataset CQL IQL TD3+BC Onestep RL RvS-R Filt. BC DT DWBC ARSQ (Ours)
halfcheetah-m 44.0 47.4 48.3 48.4 41.6 42.5 42.6 *41.4 43.7 ± 0.6
hopper-m 58.5 66.3 59.3 59.6 60.2 56.9 67.6 *56.0 99.2 ± 0.5
walker2d-m 72.5 78.3 83.7 81.8 71.7 75.0 74.0 *72.3 81.2 ± 0.9
halfcheetah-mr 45.5 44.2 44.6 38.1 38.0 40.6 36.6 38.9 41.1 ± 0.1
hopper-mr 95.0 94.7 60.9 97.5 73.5 75.9 82.7 73.0 90.7 ± 4.4
walker2d-mr 77.2 73.9 81.8 49.5 60.6 62.5 66.6 59.8 74.0 ± 2.6
halfcheetah-me 91.6 86.7 90.7 93.4 92.2 92.9 86.8 *93.1 92.4 ± 1.2
hopper-me 105.4 91.5 98.0 103.3 101.7 110.9 107.6 *110.4 110.9 ± 1.0
walker2d-me 108.8 109.6 110.1 113.0 106.0 109.0 108.1 *108.3 107.9 ± 0.3
Total 698.5 692.4 677.4 684.6 645.5 666.2 672.6 653.2 741.1
The results are presented in Tab. 1. All data is sourced from the respective papers, with the reevaluated DWBC results marked by “*”. Both ARSQ and re-evaluated DWBC are assessed using 10 trajectories over three random seeds. ARSQ demonstrates superior overall performance compared to the other baselines, indicating its capability to effectively manage suboptimal data under fully offline setting.
@@ 5.4 Ablation Studies
Ablation on Auto-regressive Conditioning.
@@
We report results on hopper-medium-expert and hopper-medium-replay from D4RL, as well as Open Oven from RLBench, all evaluated across three random seeds. As depicted in Fig. 7, Swap demonstrates a slight decline in performance, underscoring the effectiveness of the current conditioning order design. Additionally, removing any of the components degrades performance to varying degrees. When all components are removed, as in Plain, the performance is at its lowest, emphasizing the significance of dimensional and coarse-to-fine action generation.
@@
We report results on hopper-medium from D4RL and Open Oven from RLBench over three random seeds. As shown in Fig. 8, the standard ARSQ consistently performs well in both environments. In contrast, using either a level-shared or separate backbone results in diminished performance. This demonstrates the effectiveness of the shared backbone design.
@@ 6 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
@@ B.3 Network Architecture
In RLBench tasks, observations consist of a combination of RGB images and low-dimensional states. To compute the dimensional soft advantage for a given dimension, we first input the RGB images and low-dimensional states into a Convolutional Neural Network (CNN) (Li et al., 2022) encoder and a Multi-Layer Perceptron (MLP) (Haykin, 1998) encoder, respectively, to extract feature representations. These representations are then used to predict the soft value. Concurrently, the feature representations are combined with actions from previous dimensions and coarse-to-fine levels to create auto-regressive conditioning. An MLP-based shared backbone and output head are then utilized to determine the dimensional soft advantage for the given dimension.